# Measuring Value Expressions in Social Media Posts

### v2

@@ Introduction
Following this primary task, participants completed a custom developed Value Calibration Questionnaire (VCQ) of 25 additional ratings on a shared set of questions (see Table A1 and the below paragraph for a description of how these questions were sourced). Finally, participants answered a number of demographic and political questions.
We developed the Value Calibration Questionnaire as an individual difference measure to explain disagreements in value annotations. Each of the 25 questions presented a post and asked participants to rate the extent to which a specific value was reflected in the post, phrased as “To what extent does this post reflect {value}?” To ensure that these VCQ questions are highly predictive of interpersonal differences for modeling, we selected them based on a pre-study where we identified those post-value pairs that explain maximal variation in value expression annotation disagreements. In this pre-study, 51 participants rated the same set of 30 posts on all 19 values using the same procedure as this study. Unlike our other annotation process, which randomly sampled posts for annotation by each participant, here the result was a dense matrix where every participant annotated all 30 posts. We then transform the data into columns for each participant (N=51) and rows for each post/value pair ($19\times 30=570$). We demean this matrix row-wise and conduct principal component analysis (PCA) to compute bases of maximal variation (“eigenraters”). The first 25 eigenraters explain 86% of variation. For each of these eigenraters, we find the extremal post/value pair with the highest absolute value and add it to our set of questions. The full set is shown in Table A1.
To familiarize them with the task and to determine eligibility for proceeding to the main labeling task, participants first completed a short training for four synthetic posts followed by a gating task. The training introduced annotators to the meaning of the Schwartz values, while the gating task assessed their understanding. These training posts were created to unambiguously reflect particular values, and participants were not allowed to proceed until they checked the right answer. If they answered incorrectly, a message appeared giving them the correct answer and asking them to try again.
After training, they proceeded to the gating phase, where they rated four posts that, based on a pre-study, either unambiguously contained or did not contain the values of self-direction action or face. Participants who got one or fewer of the four correct were not allowed to continue. We took care in selecting the questions for both pre-tasks and in setting the eligibility threshold for the gating task to ensure a shared baseline understanding. See Figure AA1 Bottom for screenshots for the annotation of one of the gating posts.
After completing basic attention screeners and training, participants who successfully passed the gating phase proceeded to rate 30 posts drawn at random from the larger set of 5,227 posts for value expression using the branching method discussed above. When presenting a tweet to our annotators for value annotation, to convey context, the text of a given tweet (tweet_1) made in reply to or quoting another tweet (tweet_2) was represented as “tweet_1 REPLY TO: tweet_1” and “tweet_1 QUOTED: tweet_1” respectively. The annotation dataset we create is available for researchers on a case by case basis to preserve the privacy of the participants. It is stored securely in an de-identified manner.
@@
First, to improve performance beyond zero-shot GPT-4o, we perform finetuning on high-consensus value labels. To produce this training set, we randomly select 1000 posts for which we have more than 6 ratings per post from our dataset of 5,211 posts annotated by 1,079 participants across the 19 values. Then, for each post, we compute the Spearman rank correlation in value ratings for each pair of raters and average these correlation scores into an aggregate consensus score for each post. We then use the 600 highest consensus annotations of these posts (e.g., average value score rounded to the nearest integer, 11.5% of posts) as labels to fine-tune GPT-4o via the OpenAI API, producing a model that estimates the consensus label of the annotators. The fine-tuned model reached a final error of 0.0722 after 600 steps.
Then, to capture the subjective nature of value expression annotation, we combine the consensus fine-tuned model with a personalized calibration model to predict the values that a given individual would perceive as expressed in a given post. We do so by fitting a series of random forest models (one for each value) trained on the ratings of 3000 posts stratified by number of ratings to predict a given individual’s value annotation of a given post using (a) the consensus predictions from our fine-tuned model as well as (b) that individual’s responses to the 25 calibration questions. So, this model does not directly model individuals—it uses the 25 calibration questions as individual-level features. For each post–user pair, these models predict that user’s 0–6 rating for a given value V. We train 19 such models (one per value), each taking as inputs (a) the consensus model’s predictions for all values for that post and (b) the user’s responses to the 25 calibration questions.
@@
For our evaluation metric, we chose Spearman rank correlation for several reasons. For one, a correlation metric allows us to compare the relative expression across all 19 low-level Schwarz values, which follows the relational nature of the value system itself. Further, a correlation computation across values is inherently scale-invariant, allowing us to account for individual level differences in how the scale is used. We chose Spearman rank correlation because the rank ordering captures the ordinal prioritization and relational nature of values consistent with the circumplex theory. However we note that our results are qualitatively similar when using Pearson correlation instead of Spearman. In addition, a correlation metric handles the ordinal nature of the value annotations, which exist on a 7-point Likert scale.
@@ Results
We also find subjectivity in value ratings, which helps explain the sources of disagreements we identify above. To test this claim we predict individual’s value annotations across the 19 values by regressing, for each value, that particular value against each of the individual values that that individual holds, as well as demographics such as age and partisanship. The resulting regression coefficients are shown in Figure A2. We find that for 8/19 of the values, the extent to which an individual holds that value is predictive of the likelihood that they would annotate a given post with that value (dominance, resource, achievement, self-directed thoughts, self-directed actions, stimulation, personal security, and societal security), which suggests people may project their own values into the perception of those values onto novel situations. We observe some partisan differences, where Republicans are more likely to rate a given post as reflecting the values of hedonism, self-directed thoughts, self-directed actions, stimulation, humility, and nature, which may map onto cultural-political divides. Within our sample, we do not detect age effects, except for interpersonal conformity.
@@
Self-directed actions $1.118\pm 0.038$ $0.822\pm 0.025$
@@
Societal security $1.206\pm 0.036$ $0.779\pm 0.024$
Tradition $1.169\pm 0.035$ $0.777\pm 0.024$
Rule conformity $1.175\pm 0.037$ $0.737\pm 0.024$
Interpersonal conformity $1.048\pm 0.035$ $0.699\pm 0.022$
Humility $1.620\pm 0.035$ $0.971\pm 0.024$
Dependability $1.512\pm 0.037$ $0.917\pm 0.025$
Caring $1.622\pm 0.039$ $1.027\pm 0.025$
Universal concern $1.389\pm 0.039$ $0.865\pm 0.025$
Preservation of nature $0.738\pm 0.031$ $0.497\pm 0.022$
Tolerance $1.208\pm 0.036$ $0.740\pm 0.023$
@@
While fine-tuning an LLM on high-consensus labels yields a model that people agree with more than they agree with other people, the inherent subjectivity of the task underscores the point that any single model will not be performant for everyone. Therefore, next we train an interprevist model that attempts to model an individual’s subjective perception of value expression. In particular, the personalized calibration model uses a series of random forest models to predict the values that a given individual would see reflected in a given post. Using only an individual’s calibration ratings and the target post’s aggregate GPT-predicted values as inputs, the random forests achieves an average Spearman rank correlation of 0.334 with human ratings, a rate of agreement surpassing the wisdom of the crowds ($p<0.001$, dark red in Figure 4). This most performant model reflects a 66% relative improvement in correlation vs. the human-to-human baseline. This model also reflects a 28% relative improvement in correlation vs. the more difficult human-to-consensus condition. We find that for each value, the random forest models typically use the underling GPT consensus prediction of that value as the most important variable (see Figure  A3) but also make use of the calibration questions from Table A1.
@@ Discussion
In this paper, we introduce and validate a framework for measuring the expression of basic human values on social media. From a methodological perspective, we note that the large variation in value perceptions across individuals suggests that while indeed there is a degree of shared understanding, there is no singular ground truth in every case: the low levels of agreement in value annotations we observe appear even after training and gating to encourage shared understanding.
@@
To acknowledge the inherent subjectivity of the task, we introduce a specialized framework for using LLMs for annotating subjective social science constructs. With this framework, we showcase a personalized LLM-based model that people agree with more than the base model, other people, or even consensus annotations. This approach represents a perspectivist approach (fleisig2024perspectivist) to modeling subjective constructs, which we believe is critical in the age of ontologically monolithic LLMs (haghighi2025ontologies).
@@
In light of the pros and cons of the Schwartz value system, we note that the measurement framework introduced here is agnostic to the underlying value system and can be flexibly adapted to other systems with ease. Future work should explore contrasting or complementing value systems well-suited for the particular domain (e.g., the social media context). This might include other existing theories of values such as the World Values Survey or Moral Foundations Theory, or methods for deriving bottom-up theories of values via attentional probes (klingefjord2024human), or inverse reinforcement learning (oliveira2023culturally), or combinatorially combining “moral molecules” (curry2022moral). Another limitation is that this study recruited Americans to annotate X posts. How this generalizes beyond the US content and to other social media platforms remains important future work. We also note we used two GPT‑based prefilters (comprehensibility, NSFW) in addition to a GPT‑based upsampling by predicted value to identify relevant posts that conformed to our IRB. Future work should investigate how value labeling extends to incomprehensible, value-sparse and NSFW posts.
@@
There are several implications of this research that are important to keep in mind. One is the implications for feed design. If a platform understands how an individual perceives values in content, it could tune their feed to highlight posts that resonate with their values. This opens the door to new feed designs that provide increased user control and alignment. This measurement also provides a more precise way to audit what values are currently being amplified by existing feed ranking algorithms. Finally, our results underscore the need for further research into the subjectivity of social constructs, particularly when they are measured in high stakes sociotechnical contexts.
@@
Our approach imagines a world in which value expressions can be readily assessed and measured. On one hand, this might open the door to new forms of persuasion and micro targeting. On the other, it might enable new forms of ranking algorithms that rank by values. This possibility raises important questions for who gets to determine this information (platforms? users?) and how are these value profiles stored. We note that our method for personalizing value predictions and the corresponding VCQ do not rely on any privacy-compromising or sensitive individual attributes, and therefore representing a promising and procedurally legitimate approach to store value profiles.
Furthermore, there is the concern that the value-based ranking approach outlined above could induce value-based filter bubbles. Future works should explore the potential to identify and design for bridging values that are held across partisan divides.
We note that our approach relies on the OpenAI API, which if used in a real social media setting, could cause privacy concerns for users who have not explicitly consented to sharing their feeds’ content or values with a third-party service. This could be rectified with local LLM implementations. This work has been approved by the IRB of our institution to ensure adherence to ethical standards.
Acknowledgments.
We thank the HAI Social Media AI working group at Stanford, the Digital Economy Lab, Stochastic Labs, HAI, and Project Liberty. Isabel O. Gallegos is supported by the Fannie & John Hertz Foundation and Stanford Knight-Hennessy Scholars graduate fellowship. We thank Alek Chakroff and Anna Leshinskaya for helpful discussion of the Schwartz Value System and Will Schultz for helpful feedback on this paper. This project was supported by the Stanford Institute for Human-Centered Artificial Intelligence and National Science Foundation grant IIS-2403433.

### v3

@@ Methods
Following this primary task, participants completed a custom developed Value Calibration Questionnaire (VCQ) of 25 additional ratings on a shared set of questions (see Table A3 and the Personalization framework in the Methods sections for a description of how these questions were sourced). Finally, participants answered a number of demographic and political questions.
To familiarize them with the task and to determine eligibility for proceeding to the main labeling task, participants first completed a short training for four synthetic posts followed by a gating task. The training introduced annotators to the meaning of the Schwartz values, while the gating task assessed their understanding. These training posts were created to unambiguously express particular values, and participants could not proceed until they checked the right answer: if they answered incorrectly, a message appeared providing the correct answer and asking them to try again.
After training, they proceeded to the gating phase, where they rated four posts that, based on a pre-study, either unambiguously contained or did not contain the values of Self-directed action or Face. Participants who got one or fewer of the four correct were not allowed to continue. We took care in selecting the questions for both pre-tasks and in setting the eligibility threshold for the gating task to ensure a shared baseline understanding. See Figure AA3 Bottom for screenshots for the annotation of one of the gating posts.
After completing basic attention screeners and training, participants who successfully passed the gating phase proceeded to rate 30 posts drawn at random from the larger set of 5,227 posts for value expression using the branching method discussed above. When presenting a tweet to our annotators for value annotation, to convey context, the text of a given tweet (tweet_1) made in reply to or quoting another tweet (tweet_2) was represented as “tweet_2 REPLY TO: tweet_1” and “tweet_2 QUOTED: tweet_1” respectively. The annotation dataset we create is available for researchers on a case by case basis to preserve the privacy of the participants. It is stored securely in a de-identified manner.
@@
To improve performance beyond zero-shot GPT-4o, we perform finetuning on high-consensus value labels. To produce this training set, we randomly select 1000 posts for which we have more than 6 ratings per post from our dataset of 5,211 posts annotated by 1,079 participants across the 19 values. Then, for each post, we compute the Spearman rank correlation in value ratings for each pair of raters and average these correlation scores into an aggregate consensus score for each post. We then use the 600 highest consensus annotations of these posts (e.g., average value score rounded to the nearest integer, 11.5% of posts) as labels to fine-tune GPT-4o via the OpenAI API, producing a model that estimates the consensus label of the annotators. The fine-tuned model reached a final error of 0.0722 after 600 steps.
Personalization framework
To capture the subjective nature of value expression annotation, we combine the consensus fine-tuned model with a personalized calibration model to predict the values that a given individual would perceive as expressed in a given post. We do so by fitting a series of models (one for each value) trained on the ratings of 3000 posts (with 1496 heldout for evaluation) stratified by number of ratings to predict a given individual’s value annotation of a given post using (a) the consensus predictions from our fine-tuned model as well as (b) that individual’s responses to the calibration questions designed to explain variation in rater disagreement (see Figure 4). First, we describe how the calibration questions were developed, then we discuss the modeling procedure.
[…]
We de-mean this matrix row-wise and conduct principal component analysis (PCA) to compute bases of maximal variation (“eigenraters”). The first 25 eigenraters explain 86% of variation in disagreements across raters. For each of these eigenraters, we find the extremal post/value pair with the highest absolute value and add it to our set of questions, which we dub the Value Calibration Questionnaire (VCQ) to use for the main annotations study to solve the cold start problem. Each of the 25 questions presents a post and asked the participants to rate the extent to which a specific value was reflected in the post, phrased as “To what extent does this post reflect {value}?” The full set is shown in Table A3.
Given this data-driven approach to decomposing variance in value disagreements, post/value pairs are selected for inclusion because they are extremal within an eigenrater, not because they exhibit coverage of the items or values. We therefore note the appearance of both duplicate posts that sparked disagreement across different values (e.g., indices 3/16/22, 5/17, 7/18, 10/11, 2/25, and 13/23 in Table A3) and corresponding duplicate values that sparked disagreement across different posts (Caring, Tradition, Equality, Achievement, and Independent thinking). This also means certain values do not appear in the VCQ such as Power, Societal security, Respect, Humility, Responsibility, and Connection to nature, which is one limitation of this approach (see Discussion).
Our personalized framework does not directly model individuals—it instead uses the 25 calibration questions as individual-level features to solve the cold start problem (see Figure 4). For each post–user pair, these models predict that user’s 0–6 rating for a given value V using a random forest (e.g., one random forest model for each of the 19 values). Each model was trained in R using the randomForest package and default hyperparameters (500 trees). Figure A5 shows the total decrease in node impurities (variable importance) of each of the 19 random forest models, revealing how for a given value $v$ the random forest predicting $v$ primarily uses post-level the fine-tuned GPT prediction for value $v$ (shown on the top diagonal).
Figure 4: Our personalization framework uses a bespoke set of annotator features to calibrate predictions from the fine-tuned LLM.
@@
In settings where there is high levels of subjective variation exists in the labels, a traditional content-level ground truth set is not an appropriate measure and instead evaluations should rely on measure agreement with disaggregated individual labels (Gordon et al. 2022, 2021). For our evaluation metric, we chose Spearman rank correlation for several reasons. For one, a correlation metric allows us to compare the relative expression across all 19 low-level Schwartz values, which follows the relational nature of the value system itself. Further, a correlation computation across values is inherently scale-invariant, allowing us to account for individual level differences in how the scale is used. We also chose Spearman rank correlation because the rank ordering captures the ordinal prioritization and relational nature of values consistent with the circumplex theory of the Schwartz theory of Basic Human Values. However, future work might explore extensions that explicitly use the circular geometry of the circumplex directly (see Discussion). We note that our results are qualitatively similar when using Pearson correlation instead of Spearman. In addition, a correlation metric handles the ordinal nature of the value annotations, which exist on a 7-point Likert scale.
@@ Results
We also find subjectivity in value ratings, which helps explain the sources of disagreements we identify above. To test this claim we predict individual’s value annotations across the 19 values by regressing, for each value, that particular value against each of the individual values that that individual holds, as well as demographics such as age and partisanship. The resulting regression coefficients are shown in the bottom of Figure A4. We find that for 8/19 of the values, the extent to which an individual holds that value is predictive of the likelihood that they would annotate a given post with that value (Dominance, Resources, Achievement, Self-directed thoughts, Self-directed actions, Stimulation, Personal security, and Societal security), which suggests people may project their own values into the perception of those values onto novel situations. We observe some partisan differences, where Republicans are more likely to rate a given post as reflecting the values of Hedonism, Self-directed thoughts, Self-directed actions, Stimulation, Humility, and Nature, which may map onto cultural-political divides. Within our sample, we do not detect age effects, except for the value of Interpersonal conformity.
@@
Humility $1.620\pm 0.035$ $0.971\pm 0.024$
Caring $1.622\pm 0.039$ $1.027\pm 0.025$
Dependability $1.512\pm 0.037$ $0.917\pm 0.025$
Universal concern $1.389\pm 0.039$ $0.865\pm 0.025$
Hedonism $1.325\pm 0.037$ $0.927\pm 0.026$
Face $1.208\pm 0.037$ $0.918\pm 0.024$
Tolerance $1.208\pm 0.036$ $0.740\pm 0.023$
Societal security $1.206\pm 0.036$ $0.779\pm 0.024$
Rule conformity $1.175\pm 0.037$ $0.737\pm 0.024$
Tradition $1.169\pm 0.035$ $0.777\pm 0.024$
Self-directed actions $1.118\pm 0.038$ $0.822\pm 0.025$
@@
Interpersonal conformity $1.048\pm 0.035$ $0.699\pm 0.022$
Achievement $1.005\pm 0.037$ $0.749\pm 0.022$
@@
While fine-tuning an LLM on high-consensus labels yields a model that people agree with more than they agree with other people, the inherent subjectivity of the task underscores the point that any single model will not be performant for everyone. Therefore, next we train a perspectivist model that attempts to model an individual’s subjective perception of value expression. In particular, the personalized calibration model uses a series of random forest models to predict the values that a given individual would see reflected in a given post. Using only an individual’s calibration ratings and the target post’s aggregate GPT-predicted values as inputs, the random forests achieves an average Spearman rank correlation of 0.334 with human ratings, a rate of agreement surpassing the wisdom of the crowds ($p<0.001$, dark red in Figure 4). This most performant model reflects a 66% relative improvement in correlation vs. the human-to-human baseline. This model also reflects a 28% relative improvement in correlation vs. the more difficult human-to-consensus condition. We find that for each value, the random forest models typically use the underling GPT consensus prediction of that value as the most important variable (see Figure  A5) but also make use of the calibration questions from Table A3.
@@
What Drives Personalization Improvements?
Adjusting differentially subjective values
In Table 1, we see that the most overall subjective values in our sample were humility, caring, and dependability, while some of the most agreed upon values were resources, dominance, and personal security. To understand which values are being adjusted and how that effects personalization improvements, for a given value $v$, we compute the average absolute difference ($\Delta$) between the predictions of the fine-tuned consensus and the personalized models for value $v$. We find that the top adjusting values are humility ($\Delta=0.771$), caring ($\Delta=0.655$) and dependability ($\Delta=0.655$), and the least adjusting values are personal security ($\delta=0.344$) and preservation of nature ($\Delta=0.429$) with an overall value-wise spearman rank correlation of R=0.833. This indicates that personalization improves agreement by differentially adjusting subjective values where perceptions differ.
Personalized to whom?
For each participant, we compute the average spearman rank correlation between that participant’s ratings of a given post and the model predictions and then compute the difference in average spearman rank correlation between the consensus model and the personalized. In other words, we compute a participant-level measure of the delta between the red bar in Figure 5 and the dark red bar in Figure 5. We then use a linear regression to compute the association between this delta (e.g. performance gain due to personalization) with individual differences. We find that personalization gains are relatively higher for men ($p=0.0102$), Republicans ($p=0.0297$) and no effect across ages ($p=0.824$).
When comparing posts sourced from the For You page versus the Following page, we see no differences in either personalization gains ($p=0.709$) or individuals average level of disagreement with the consensus annotation ($p=0.8$).
@@ Discussion
In this paper, we introduce and validate a framework for measuring the expression of basic human values on social media paradigmatic of a shift towards perspectivism. From a methodological perspective, we note that the large variation in value perceptions across individuals suggests that while indeed there is a degree of shared understanding, there is no singular ground truth in every case: the low levels of agreement in value annotations we observe appear even after training and gating to encourage shared understanding.
@@
To acknowledge the inherent subjectivity of the task, we introduce a specialized framework for using LLMs for annotating subjective social science constructs. With this framework, we showcase a personalized LLM-based model that people agree with more than the base model, other people, or even consensus annotations. This approach represents a perspectivist approach (Fleisig et al. 2024) to modeling subjective constructs, which we believe is critical in the age of ontologically monolithic LLMs (Haghighi et al. 2025).
We note that in this paper, we were primarily focused on arguing for a paradigm shift towards personalization, and our personalization framework using random forests and a spectral decomposition of variations in disagreement is intended to demonstrate one possible instantiation that yields improvements beyond consensus annotations.
The personalization framework presented here therefore represents one exemplary approach as a proof-of-concept but future work is needed to evaluate this approach more deeply and contrast it to a more comprehensive design space. For example, as discussed above, the method of identifying highly predictive features to solve the cold start problem does not include certain values and thus requires future work to understand how items identified generalize from one sample of raters (and items) to another and develop alternative methods for solving the cold start problem efficiently.
@@
In addition, we note that we only considered perspectivist models for the Schwartz value system. Future work should explore adapting this approach to other value systems that may be well-suited for the particular domain (such as the social media context). Depending on the domain, this might include other existing theories of values such as the World Values Survey or Moral Foundations Theory, or methods for deriving bottom-up theories of values via attentional probes (Klingefjord et al. 2024), or inverse reinforcement learning (Oliveira et al. 2023), or combinatorially combining “moral molecules” (Curry et al. 2022). In addition, we note that the metric we used, Spearman rank correlaiton, does not explicitley capture the circualar geometry of the Shwartz value system and future work could explore bespoke measures that capture angular distance more precisely.
Another limitation is that this study recruited Americans to annotate X posts. How this generalizes beyond the US content and to other social media platforms remains important future work. We also note we used two preliminary GPT‑based filters (comprehensibility, NSFW) in addition to a GPT‑based upsampling by predicted value to identify relevant posts that conformed to our IRB. Future work should investigate how value labeling extends to incomprehensible, value-sparse and NSFW posts.
There are several both epistemic and practical implications of this shift towards personalized values that are important to keep in mind. Epistemically, this underscores the need for further research into the subjectivity of a variety of social constructs (such as “helpfulness” or “toxicity”), particularly when they are measured and deployed in high stakes sociotechnical contexts. These constructs may operate differently for different people, and therefore these systems may inadvertently exhibit bias based on divergent interpretations.
@@
One practical implication of this work is for social media feed design. If a platform can carefully understand how a user perceives values in content, it could tune that user’s feed to highlight posts that resonate with their values. This opens the door to new feed algorithms that provide increased user control and alignment. In addition, this shift also provides a more precise way to audit what values are currently being amplified by existing feed ranking algorithms.
@@
Our approach imagines a world in which value expressions can be readily assessed and measured. This may open the door to new forms of persuasion and microtargeting based on value-based rhetoric. This could be mitigated by mandated disclosure for advertisers and opt-out mechanisms for consumers. It could also include labels for value-laden language, to help consumers identify and contextualize value aligned or value misaligned content.
This may also enable new forms of ranking algorithms that rank by values, which represents a departure from the engagement-based ranking paradigm of existing platforms. This possibility raises important questions for who gets to determine this information (platforms? users?) and how are these value profiles stored. We note that our method for personalizing value predictions and the corresponding VCQ do not rely on any privacy-compromising or sensitive individual attributes, and therefore representing a promising and procedurally legitimate approach to store value profiles.
Furthermore, there is the concern that the value-based ranking approach outlined above could induce value-based filter bubbles. This could be mitigated by identifying and design for bridging values that are held across partisan divides. However future work is needed to explore how what these bridging values are and how effective they are in mitigating value-based filter bubbles.
We note that our approach relies on the OpenAI API which, if used in a real social media setting, could cause privacy concerns for users who have not explicitly consented to sharing their feeds’ content or values with a third-party service. This could be mitigated by local LLM implementations that do not share data with a third-party, or by sufficiently explicit data-sharing agreements with users.
This work has been approved by the IRB of our institution to ensure adherence to ethical standards. In all the reproductions of social media posts from our dataset in this paper, we have reworded the post, blurred meta-data and replaced meta-data to preserve anonymity while maintaining value saliency.
@@
Appendix A Evaluating LLM’s for data processing
NSFW Classifier
To evaluate our classifier for the comparatively straightforward, unambiguous and sparse task of identifying content that is not safe for work (NSFW), we compute agreement on a random sample of 100 positive and 100 negative examples tagged by the classifier. Two authors (A1, A2) annotated the 200 posts with a binary label for NSFW based on the codebook shown in Table A5. A1 and A2 agree with each other with Cohen’s $\kappa=0.65$. A1 agreed with the model with Cohen’s $\kappa=0.70$ corresponding to an accuracy of 84.9%. A2 agreed with the model with Cohen’s $\kappa=0.86$ corresponding to an accuracy of 93.0%.
Comprehensibility Classifier
Next, we evaluate our classifier for the relatively ambiguous, context-based and label balanced task of identifying content that is comprehensible and therefore capable of being judged for values. Comprehensibility is a highly subjective task because annotators are coming from different baselines, where what is comprehensible to them may differ based on their backgrounds, experiences and contexts. To evaluate this model, we also compute agreement on a random sample of 100 positive and 100 negative examples tagged by the classifier. Three authors (A1, A2, A3) annotated the 200 posts with a binary label for comprehensibility based on the codebook shown in Table A4. We then computed an overall majority label from the three rater’s responses (121 labeled as comprehensible, 79 labeled as incomprehensible) and downsample the comprehensible posts to balance the labels with respect to the human-rated majority labels.
For this balanced set of annotations, the classifier has a recall of 0.695 and a precision of 0.721 (see Table A1). However many of the misclassifications occur for posts where raters disagreed on if the post was comprehensible: when considering the subset of ratings where all three raters unanimously agreed on the label (85 labeled as comprehensible, 48 labeled as incomprehensible, similarly downsampling the comprehensible posts to balance the labels with respect to the ground truth), the classifier has a recall of 0.709 and a precision of 0.813.
Table A1: Evaluation metrics for comprehensibility classifier grounded in 1) majority annotations from the three raters, and 2) majority annotations where the raters unanimously agreed.
Metric Majority Unanimous Majority
Accuracy 0.703 0.740
$\kappa$ 0.405 0.479
Recall 0.695 0.709
Precision 0.722 0.813
F1 Score 0.708 0.757
Number of posts 158 96
Preliminary Values Classifier
To evaluate the preliminary values classifier we used to upsample infrequent values into the final annotation dataset, we leverage the fact that our annotated dataset contains examples of positive and negative labels of the preliminary values classifier for each of the 19 values. In particular, for each value $v$, we compare the mean human rating for posts we sampled as positive examples (i.e. the classifier identified as containing the value $\hat{y}_{v}\geq 4$) to the other posts sampled for other values (i.e. $\hat{y}_{v}<4$), the results of which are shown in Table A2. We see that for 18/19 of the values, the average rating of the posts sampled via the preliminary values classifier is statistically greater than the those that were not.
Table A2: Schwartz Value Rating Effects
Value Avg. Rating Avg. Rating $p$-val
$(\hat{y}_{v}\geq 4)$ $(\hat{y}_{v}<4)$
Face 1.174 0.761 <0.001<0.001
Dominance 0.472 0.371 0.001
Resources 0.844 0.390 <0.001<0.001
Achievement 1.245 0.517 <0.001<0.001
Hedonism 2.457 0.955 <0.001<0.001
Self-Dir. Thoughts 1.423 0.719 <0.001<0.001
Self-Dir.Actions 1.322 0.734 <0.001<0.001
Stimulation 1.238 0.583 <0.001<0.001
Personal Security 0.627 0.208 <0.001<0.001
Societal Security 1.389 0.649 <0.001<0.001
Tradition 1.677 0.683 <0.001<0.001
Rule Conformity 1.989 0.620 <0.001<0.001
Interp. Conformity 0.787 0.631 <0.001<0.001
Humility 1.426 1.345 0.171
Dependability 1.865 1.235 <0.001<0.001
Caring 2.551 1.235 <0.001<0.001
Universal Concern 2.446 0.851 <0.001<0.001
Pres. of Nature 3.076 0.465 <0.001<0.001
Tolerance 1.803 0.754 <0.001<0.001
Appendix B Case studies of value disagreements in social media posts
What kind of disagreements did we notice in social media posts? While we did not observe any global linguistic or contextual trends associated disagreements, we did observe many disagreements where certain value dimensions of a post where salient to some participants, while others dimensions were salient to others. In addition to annotating the posts for values, our survey also provided participants an optional space to provide their rationale. These rationales help shed light on some disagreements and divergent perspectives within the posts.
Figure A1: Anonymized example post from our dataset.
Case Study 1:
Figure A1 depicts one post from our dataset, describing a magnetic storm hitting earth. On one hand, One participant read the post as a socially meaningful act—inferring care or universal concern from the decision to share timely information about a potentially impactful event, stating “great post to give others information about world shifting weather events.” On the other hand, another participant grounds value inference in the described phenomenon itself, focusing on the natural event’s aesthetic and reminder of humans’ smallness in the face of nature, mapping it onto humility and appreciation of nature, stating “The northern lights are just beautiful and a good reminder of the beauty of earth.”
Case Study 2:
Figure A2 depicts another post from our dataset, describing an NFL playing giving a shout out to their 4th grade teacher. One participant centers the relational and moral dimension of the story, interpreting the public acknowledgment of a teacher as an expression of dependability and caring toward those who supported the athlete’s development, stating: “it’s important to thank those who inspire us”. The other focuses on the athlete’s success itself, interpreting the post as a story about achievement and public recognition, stating: “It is a positive story about the guy who was drafted. It is about his success.”
Figure A2: Anonymized example post from our dataset.
Appendix C Attention checks